OpenStudy (anonymous):

what is the value of M & N respectively? if M39048458N is divisible by 8 & 11; where M & N are single digit integer.

OpenStudy (anonymous):

For a number to be divisible by 8, we check if the number formed by its last three digits (hundreds, tens and ones) is divisible by 8 or not so we check for what value of N is 58N divisible by 8 so we start by checking for 580 580 divided by 8 = 72.5 so logically 73 x 8 = 584 is divisible by 8 so N = 4

OpenStudy (anonymous):

for division by 8, value of M can be anything from 1 to 9

OpenStudy (anonymous):

but the second condition is also given that the no should be divisible by 11 also

OpenStudy (anonymous):

for M39048458N to be divisible by 11, the difference of the sums of alternate digits should be either zero or a multiple of 11 taking sum of alternate digits (starting from the ones digit) since we already know that N=4 SUM I = N+5+8+0+3 = 4+16 = 20 SUM II = 8+4+4+9+M = M+25 difference SUM II - SUM I = (M+25) - 20 = M+25-20 = M+5 if difference = 0, then M+5 = 0 or M = -5 which is not possible if difference is 11, then M+5 = 11 or M = 6 which is workable so M = 6 and N=4

OpenStudy (anonymous):

hope now it is complete for you....... Pls give medal if u r satisfied.....

OpenStudy (anonymous):

good answer

OpenStudy (anonymous):

u r welcome☺

OpenStudy (anonymous):

How many numbers less than1000 which have more than 24 factors?

OpenStudy (anonymous):

ise bhi try karna

OpenStudy (anonymous):

is this a school question or r u posing a puzzle?

OpenStudy (anonymous):

these questions r related to CAT exam